6. (a) Draw the logic circuit for the following expression: AB + A(B+C)

(b) Simplify the expression in 6(a) by using the rules of Boolean algebra provided in (5). (c) Draw the simplified logic gate circuit derived in (b)



1
Expert's answer
2021-11-24T05:44:10-0500

(a) Let us draw the logic circuit for the expression "AB + A(B+C):"



(b) Let us simplify the expression "AB + A(B+C)" by using the rules of Boolean algebra:


"AB + A(B+C)=A(B+(B+C))=A((B+B)+C)=A(B+C)"


(c) Let us draw the simplified logic gate circuit derived in (b):